OpenStudy (anonymous):

with medal Help Plz!!!!!!!! What is the sum of the geometric sequence 1, -6, 36, ... if there are 6 terms?

OpenStudy (anonymous):

If the first term of a geometric sequence is \(\Large a\) and the common ratio is \(\Large r\), then: -the second term is \(\Large ar\); -the third term is \(\Large ar^2\); etc. In general, the nth term is \(\Large ar^{n−1}\) If we sum the first \(\Large n\) terms of a geometric series, we obtain: \(\Large S_n=a+ar+....+ar^{n−2}+ar^{n−1}\) So,the geometric sequence 1, -6, 36,... has \(\Large a=1, r=-6,n=6\) the sum:\(\Large S=1-6+36-216+1296-7776=-6665\)